# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ 5.1 Constant LRS
For constant LRS, we have the following result, with the proof deferred to Appendix D.1.
@@
Under Assumption 5.1, we have
$\mathcal{E}_{K}\eqsim M^{-s\beta}+(\eta K)^{-s}+\frac{\eta}{B}\left(\sigma^{2}+(\eta K)^{-(2-\frac{1}{\beta})}\right).$
In this case, the noise term is simplified to $\frac{\eta}{B}\big(\sigma^{2}+(\eta K)^{-(2-\frac{1}{\beta})}\big)$. The fit-dependent component $(\eta K)^{-(2-\frac{1}{\beta})}$ is indeed dominated by the full-batch GD term $(\eta K)^{-s}$ when $s\leqslant 2-\frac{1}{\beta}$. In the subsequent statement, we shall omit this term for clarity. Let $\gamma:=\eta/B$ be the effective learning rate. Then, the scaling law can be rewritten as
@@
where the excess risk depends the learning rate via $\gamma=\eta/B$. This suggests that we should scale the learning rate linearly with respect to batch size (a.k.a. linear scaling rule) (krizhevsky2014one, ; goyal2017accurate, ; mccandlish2018empirical, ).
@@
We observe that both the best achievable excess risk $\mathcal{E}_{\mathrm{opt}}$ and optimal learning rate $\gamma_{\mathrm{opt}}$ depend exclusively on the task’s relative difficulty $s$. For a fixed target (fixed $\alpha$), a higher‑capacity model (smaller $\beta$) gives a larger $s=\alpha/\beta$ and is therefore more data‑efficient.
@@
$\gamma_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{1+(s+1)\beta}},\,\,M_{\mathrm{opt}}\eqsim C^{\frac{1}{1+(s+1)\beta}},\,\,D_{\mathrm{opt}}\eqsim C^{\frac{(s+1)\beta}{1+(s+1)\beta}},\qquad\mathcal{E}_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{1+s\beta+{\beta}}}.$
This shows that the performance of the compute-optimal model improves with the budget $C$ in a power law. For a fixed task ($\alpha=s\beta$ fixed), increasing the model capacity ( $\beta\downarrow$) can enhance the compute efficiency – the extra $\beta$ in the scaling exponent $\frac{s\beta}{1+s\beta+{\beta}}$ quantifies this gain. This provides an explanation of practical observations in LLM pre-training: Large models are more compute efficient than small models kaplan2020scaling ; hoffmann2022training . In addition, the optimal learning rate $\gamma_{\mathrm{opt}}$ decreases as $C$ grows and the optimal strategy allocates compute more aggressively to data than to model – again consistent with current LLM pre-training practices (bi2024deepseek, ; shuai2024scaling, ; hoffmann2022training, ).
bordelon2024feature also studied the compute-optimal scaling for a similar setting but assumed a fixed learning rate and an absence of label noise. In contrast, we consider a more realistic scenario where the learning rate is optimally tuned and the irreducible risk is present. This enables us to derive a compute-optimal scaling that matches empirical observations.
@@ 5.2 Exponential Decay LRS
We next consider the exponential decay (exp-decay) LRS for a given number of training steps $K$ (ge2019step, ; wu2022last, ):
$\varphi(\tau)=a\exp(-\lambda\tau),\qquad\varphi(Kh)=b,$
where $\lambda$ is chosen such that $\varphi(Kh)=b$. For brevity, we assume $h=1$. Note that the two hyperparameters $a$ and $b$ denote the maximal and minimum learning rates, respectively.
@@
$\mathcal{E}_{K}\eqsim M^{-s\beta}+T^{-s}+\sigma^{2}\left(\frac{b}{B}+(a-b)\frac{\min\{M,T^{1/\beta}\}}{BT}\right),$
where $T=(a-b)K/\log(a/b)$ is the total intrinsic training time.
Let $a=b/K$. Then, the intrinsic time becomes $T=a(K-1)/\log K$. In contrast, a constant LRS with $\eta=a$ achieves $T=aK$. Thus, exp-decay LRS incurs only a $\log K$ reduction in the intrinsic time, even if it drives the learning rate below $a/K$. Let $\gamma=a/B$ be the maximal effective learning rate.
Data-optimal scaling. In this regime, we have $M_{\mathrm{opt}}=\infty$ and:
  • If $s\geqslant 1-1/\beta$ (easy-learning), then $\gamma_{\mathrm{opt}}\eqsim(D/\log D)^{-\frac{1+s\beta-\beta}{1+s\beta}}$ and $\mathcal{E}_{\mathrm{opt}}\eqsim(D/\log D)^{-\frac{s\beta}{s\beta+1}}$.
  • If $s<1-1/\beta$ (hard-learning), then $\gamma_{\mathrm{opt}}\eqsim 1$ and $\mathcal{E}_{\mathrm{opt}}\eqsim(D/\log D)^{-s}$.
The derivation can be found in Appendix D.2. The hard-learning regime arises due to the stability constraint: learning rates must remain bounded when scaling $D$ up. Relative to the constant LRS (see Eq. (8)), exp-decay LRS yields strictly faster excess‐risk decay. This provides a formal justification for the importance of learning‐rate decay in stochastic optimization.
@@
In the easy‑learning regime, we again observe that compute-optimal allocation consistently favors data over model. Notably, the optimal compute split depends solely on the task’s intrinsic difficulty and the relative fraction of data over model: $D_{\mathrm{opt}}/M_{\mathrm{opt}}\;\eqsim\;C^{\alpha/(2+\alpha)}$ decreases as the task becomes harder. The optimal $\gamma_{\mathrm{opt}}$ also decays with the compute budget $C$, and for fixed $\alpha$, higher-capacity models ($\beta\downarrow$) require smaller $\gamma_{\mathrm{opt}}$. Moreover, the excess‐risk rate is governed solely by $\alpha=s\beta$, so increasing capacity alone does not yield asymptotic gains in this easy-learning regime.
In the hard-learning regime, data still dominates compute allocation, but now the optimal split depends only on the model capacity $\beta$, independent of the relative task difficulty. Moreover, the optimal maximal learning rate remains constant ($\gamma_{\mathrm{opt}}\eqsim 1$). These results imply that a single, universal choice of compute split and learning rate suffices to attain optimal scaling across all tasks satisfying $s<1-1/\beta$, greatly simplifying hyperparameter tuning. Finally, in this regime, higher‐capacity models become strictly more compute‐efficient, as evidenced by the excess-risk scaling exponent $s\beta/(1+{\beta})$.
@@ 5.3 WSD-like LRS
We lastly turn to consider a WSD-like LRS (zhai2022scaling, ; hu2024minicpm, ), which comprises a $K_{1}$-step stable phase followed by a $K_{2}$-step decay phase, for a total $K=K_{1}+K_{2}$ steps, given by
$\varphi(\tau)=\begin{cases}a&,\text{ if }\tau\leqslant K_{1}h;\\ a\exp(-\lambda(\tau-K_{1}h))&,\text{ if }\tau>K_{1}h.\end{cases}$ (9)
where $\lambda$ is chosen such that $\varphi(Kh)=b$. For brevity, we assume $h=1$ and let $r=K_{2}/K$. This schedule is thus characterized by three hyperparameters: the largest learning rate $a$, the final learning rate $b$, and the decay proportion $r$ controlling the decay-phase length. (The warmup phase is omitted, as it does not affect our analysis.)
@@
$\mathcal{E}_{K}\eqsim M^{-s\beta}+(T_{1}+T_{2})^{-s}+\sigma^{2}\left(\frac{b}{B}+(a-b)\frac{\min\{M,T_{2}^{1/\beta}\}}{BT_{2}}\right),$
where $T_{1}=aK_{1}$ and $T_{2}=(a-b)K_{2}/\log(a/b)$denote the intrinsic training times of the stable and decay phases, respectively.
We can see that WSD-like LRS can leverage the initial stable phase to boost the intrinsic training time. In particular, for any decay‐phase fraction $r<1$, we have $T=T_{1}+T_{2}\geqslant(1-r)Ka$, which far exceeds the the intrinsic time $T\eqsim aK/\log K$ achieved by the pure exp-decay LRS. Consequently, WSD removes logarithmic factors in the full-batch GD term, without altering the noise term’s order as long as $r>0$. Building on this insights, we show that WSD can indeed improve the scaling efficiency, as detailed below.
@@
  • If $s\geqslant 1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim D^{-\frac{1+s\beta-\beta}{1+s\beta}}(\log D)^{\frac{\beta-1}{1+s\beta}}$, $r_{\mathrm{opt}}\in(0,1)$ and $\mathcal{E}_{\mathrm{opt}}\eqsim D^{-\frac{s\beta}{s\beta+1}}(\log D)^{\frac{s\beta-s}{1+s\beta}}$.
  • If $s<1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim 1$, $r_{\mathrm{opt}}\gtrsim D^{-\frac{\beta-1-s\beta}{\beta-1}}\log D$ and $\mathcal{E}_{\mathrm{opt}}\eqsim D^{-s}$.
Compared with the exp‐decay LRS, both regimes enjoy an additional logarithmic improvement in excess‐risk decay. Particularly, in the hard‐learning regime, the logarithmic factor is completely eliminated. To achieve this, the decay-phase duration only needs to scale sublinearly with $D$, as indicated by $r_{\mathrm{opt}}\to 0$ as $D\to\infty$. This observation aligns with the WSD practice in LLM pre-training – where the decay phase typically occupies only 10% of training. Moreover, our theory suggests that for harder tasks, the decay fraction could be reduced even further to enhance efficiency.
Analogous improvements hold in the compute-limited regime. Assuming $b=a/K$ and imposing the compute constraint $MD=C$, the compute-optimal scaling satisfies:
  • If $s\geqslant 1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim(C/\log C)^{-\frac{1+s\beta-\beta}{2+s\beta}},r_{\mathrm{opt}}\in(0,1),M_{\mathrm{opt}}\eqsim(C/\log C)^{\frac{1}{2+s\beta}}$, $D_{\mathrm{opt}}\eqsim C^{\frac{1+s\beta}{2+s\beta}}(\log C)^{\frac{1}{2+s\beta}},$ and $\mathcal{E}_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{2+s\beta}}(\log C)^{\frac{s\beta-s}{2+s\beta}}.$
@@ 6 Experiments
6.1 Teacher–Student Kernel Regression
We begin by empirically assessing how LRS affects the scaling behavior of kernel‑regression, thereby validating our theoretical predictions. The three – constant, exp-decay, and WSD-like – LRSs analyzed in Section 5 are evaluated on both easy ($s\geqslant 1-1/\beta$) and hard ($s<1-1/\beta$) tasks. To keep the experiments lightweight, we confine our study to the data‑limited regime; full implementation details are provided in the Appendix B.2.
Figure 2(a) compares the scaling laws obtained by numerical computation using the FSL (7) and the ones predicted by our analytical analyses in Section 5. We see that the scaling curves exhibit strong agreement across both easy and hard learning regimes for all three tested LRSs.
Figure 2(b) repeats the comparison using discrete SGD for a hard-learning task. Each configuration is run for $200$ independent trials; we plot the mean risk together with the standard deviation. Figure 2(b)(left) confirms that the empirical scaling laws for SGD closely match the theoretical predictions. Among these LRSs, WSD indeed achieves the lowest risk, followed by exponential decay, with the constant schedule performing the worst. Figure 2(b)(right) further compares the risk trajectories of SGD under the three LRSs. While WSD decays much more slowly than exp-decay during the stable phase, it initiates a sharper drop once the decay phase begins, ultimately converging to a lower final risk. This counterintuitive pattern mirrors observations in LLM pre‑training hu2024minicpm ; wen2024understanding .
Figure 2: The impact of LRS on the scaling behavior of teacher–student kernel regression. In all panels, dashed curves represent the theoretical predictions from Section 5. (a) Solid curves show the risk predicted via numerical simulation using the FSL, where the maximal learning rate is scaled as $\eta_{\max}=0.05D^{-r}$, where the exponent $r$ follows the theory in Section 5 for each LRS. (b) Discrete‑SGD experiments: (left) empirical scaling laws; (right) risk trajectories under the three LRSs, each with well‑tuned $\eta_{\max}$.
@@ 6.2 LLM Pre-training
This subsection assesses the practical relevance of FSL – originally derived for teacher–student kernel regression – as a surrogate model for capturing the loss dynamics of LLM pre‑training.
FSL can fit and predict loss curves.
We first quantify the descriptive and predictive accuracy of FSL. Following the protocol of tissue2024scaling and luo2024a , we restrict attention to the post‑warmup portion of the training trajectory. Two LLaMA (touvron2023llama, ) models (400 M and 1 B) are trained on 20 B tokens under three LRSs: cosine, WSD, and the 8‑1‑1 (bi2024deepseek, ). In the 8‑1‑1 LRS, the learning rate is reduced by a factor $\sqrt{10}$ at 80% and 90% of the total token budget, yielding a final value that is $0.1$ times the peak rate.
For each model we (i) fit the FSL parameters on the loss curve obtained using the 8‑1‑1 LRS and (ii) deploy the fitted FSL to predict the loss dynamics for the cosine and WSD schedules. Figure 3(a) demonstrates that FSL not only fits the 8‑1‑1 trajectory accurately but also generalizes reliably to the unseen WSD and cosine LRSs, across both model sizes.
@@
The FSL‑optimal LRS.
We next ask whether the fitted FSL can be leveraged to design better LRSs. To this end, we numerically minimize the terminal risk over the space of LRSs using the fitted FSL (see Appendix B.3 for details). In this experiment, we consider a 1B-parameter mixture-of-experts QwenMoE model (yang2024qwen25, ), trained on 20B tokens using the three previously discussed LRSs; see Figure 3(b) (left). We fit the FSL using the trajectory from the 8-1-1 LRS and numerically solve for the FSL-optimal LRS. The model is then trained under this FSL-optimal schedule, using the same compute budget, and compared against the baseline LRSs. The results are summarized in Figure 3(b).
Figure 3(b)(left) shows that surprisingly, the FSL-optimal LRS is WSD-like and the decay phase drives the learning rate far below the conventional $0.1\eta_{\max}$ threshold. This mirrors recent empirical recommendations by bergsma2025straight ; hagele2024scaling . Furthermore, Figure 3(b)(right) demonstrates that the FSL‑optimal schedule yields a lower terminal loss than all baseline LRSs, substantiating its practical relevance.
Taken together, these results suggest that FSL offers a faithful analytical surrogate model for studying LLM training dynamics and holds promise as a principled tool for interpreting and designing LRSs in large-scale pre-training.

### v3

@@ 5.1 Constant LRS
Under Assumption 5.1, for any $\eta K\gtrsim 1$, we have
$\mathcal{E}_{K}\eqsim M^{-s\beta}+(\eta K)^{-s}+\frac{\eta}{B}\sigma^{2}.$ (21)
Let $\gamma:=\eta/B$ be the effective learning rate. Then, the scaling law can be rewritten as
@@
where the excess risk depends the learning rate via $\gamma=\eta/B$. This suggests that we should scale the learning rate linearly with respect to batch size (a.k.a. linear scaling rule) [28, 17, 42].
@@
Notably, both the best achievable excess risk $\mathcal{E}_{\mathrm{opt}}$ and optimal learning rate $\gamma_{\mathrm{opt}}$ depend exclusively on the task’s relative difficulty $s$. For a fixed target (fixed $\alpha$), a higher‑capacity model (smaller $\beta$) gives a larger $s=\alpha/\beta$ and is therefore more data‑efficient.
@@
$\gamma_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{1+(s+1)\beta}},\,\,M_{\mathrm{opt}}\eqsim C^{\frac{1}{1+(s+1)\beta}},\,\,D_{\mathrm{opt}}\eqsim C^{\frac{(s+1)\beta}{1+(s+1)\beta}},\qquad\mathcal{E}_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{1+s\beta+{\beta}}}.$ (23)
This shows that the performance of the compute-optimal model improves with the total compute budget $C$ in a power law. For a fixed task ($\alpha=s\beta$ fixed), we have the following observations:
  • Increasing model capacity ( $\beta\downarrow$) enhances compute efficiency—the extra $\beta$ in the scaling exponent $\frac{s\beta}{1+s\beta+{\beta}}$ quantifies this gain. This explains a well-known empirical observation in LLM pre-training: Large models are more compute-efficient than small models [27, 21].
  • The optimal learning rate $\gamma_{\mathrm{opt}}$ decreases as $C$ grows, and the compute-optimal allocation favors investing more in data than in model size—again consistent with current LLM pre-training practice [5, 56, 21].
Note that [9] also investigated compute-optimal scaling for constant LRS but assumed a fixed learning rate and no label noise. In contrast, we consider a more realistic scenario where the learning rate is optimally tuned and the irreducible risk is present, leading to a compute-optimal scaling law that matches empirical observations.
@@ 5.2 Exponential Decay LRS
For a given number of training steps $K$ [16, 66], an exponential decay (exp-decay) LRS is given by
$\varphi(\tau)=a\exp(-\zeta\tau),\qquad\varphi(Kh)=b,$
where $\zeta$ is chosen such that $\varphi(Kh)=b$. For brevity, we assume $h=1$. Note that the two hyperparameters $a$ and $b$ specify the peak and final learning rates, respectively.
@@
$\mathcal{E}_{K}\eqsim M^{-s\beta}+T^{-s}+\sigma^{2}\left(\frac{b}{B}+(a-b)\frac{\min\{M,T^{1/\beta}\}}{BT}\right),$ (24)
where $T=(a-b)K/\log(a/b)\gtrsim 1$ is the total intrinsic training time.
Let $b=a/K$. Then the intrinsic time becomes $T=a(K-1)/\log K$, whereas a constant LRS with step size $\eta=a$ yields $T=aK$. Thus, exp-decay LRS drives the learning rate down to as small as $a/K$, yet sacrifices only a logarithmic factor of intrinsic time compared to the constant schedule.
Data-optimal scaling. Let $\gamma=a/B$ be the effective peak learning rate. By minimizing the right hand side of (24) with respect to $a,b,K,B,M$ under the constraint $KB=D$ (see Appendix D.2), We obtain $M_{\mathrm{opt}}=\infty$ and
  • If $s\geqslant 1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim(D/\log D)^{-\frac{1+s\beta-\beta}{1+s\beta}}$ and $\mathcal{E}_{\mathrm{opt}}\eqsim(D/\log D)^{-\frac{s\beta}{s\beta+1}}$.
  • If $s<1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim 1$ and $\mathcal{E}_{\mathrm{opt}}\eqsim(D/\log D)^{-s}$.
Compared with the constant LRS (see Eq. (22)), exp-decay LRS achieves a strictly faster decay of the excess risk, justifying the importance of learning‐rate decay in stochastic optimization.
@@
In the easy-learning regime, the excess-risk rate is determined solely by the intrinsic difficulty $\alpha=s\beta$; hence, increasing model capacity alone does not lead to asymptotic gains. The compute-optimal allocation consistently favors data over model and moreover, the optimal compute split depends solely on the task’s intrinsic difficulty, with ratio $D_{\mathrm{opt}}/M_{\mathrm{opt}}\;\eqsim\;C^{\alpha/(2+\alpha)}$ decreasing as the task becomes harder. This implies that, for harder tasks, one should allocate more compute to increasing model size. The optimal $\gamma_{\mathrm{opt}}$ decreases with the compute budget $C$, and for fixed $\alpha$, higher-capacity models ($\beta\downarrow$) require smaller $\gamma_{\mathrm{opt}}$.
In the hard-learning regime, data still dominates compute allocation, but now the optimal split depends only on model capacity, independent of the task difficulty. Moreover, the optimal maximal learning rate remains constant ($\gamma_{\mathrm{opt}}\eqsim 1$). These results imply that a single, universal choice of compute split and learning rate suffices to attain optimal scaling across all tasks satisfying $s<1-1/\beta$, greatly simplifying hyperparameter tuning. Finally, in this regime, higher‐capacity models (smaller $\beta$) become strictly more compute‐efficient, as evidenced by the excess-risk scaling exponent $s\beta/(1+{\beta})$.
@@ 5.3 WSD-like LRS
We lastly turn to consider a WSD-like LRS [71, 22], which comprises a $K_{1}$-step stable phase followed by a $K_{2}$-step decay phase, for a total $K=K_{1}+K_{2}$ steps, given by
$\varphi(\tau)=\begin{cases}a&,\text{ if }\tau\leqslant K_{1}h;\\ a\exp(-\zeta(\tau-K_{1}h))&,\text{ if }\tau>K_{1}h.\end{cases}$ (25)
where $\zeta$ is chosen such that $\varphi(Kh)=b$. For brevity, we assume $h=1$ and let $r=K_{2}/K$. This schedule is thus characterized by three hyperparameters: the peak learning rate $a$, the final learning rate $b$, and the decay proportion $r$, which controls the duration of decay-phase. (The warmup phase is omitted, as it does not affect our analysis.)
@@
$\mathcal{E}_{K}\eqsim M^{-s\beta}+(T_{1}+T_{2})^{-s}+\sigma^{2}\left(\frac{b}{B}+(a-b)\frac{\min\{M,T_{2}^{1/\beta}\}}{BT_{2}}\right),$ (26)
where $T_{1}=aK_{1}\gtrsim 1$ and $T_{2}=(a-b)K_{2}/\log(a/b)\gtrsim 1$ denote the intrinsic training times of the stable and decay phases, respectively.
We see that WSD-like LRS can leverage the initial stable phase to boost the intrinsic training time. For a decay proportion $r<1$, we have $T=T_{1}+T_{2}\geqslant(1-r)Ka$, which far exceeds the the intrinsic time $T\eqsim aK/\log K$ achieved by the pure exp-decay LRS. Consequently, WSD removes logarithmic factors in the full-batch GD term, without altering the noise term’s order as long as $r>0$. Building on this insights, we show that WSD can indeed improve the scaling efficiency, as detailed below.
@@
  • If $s\geqslant 1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim D^{-\frac{1+s\beta-\beta}{1+s\beta}}(\log D)^{\frac{\beta-1}{1+s\beta}}$, $r_{\mathrm{opt}}\eqsim 1$ and $\mathcal{E}_{\mathrm{opt}}\eqsim D^{-\frac{s\beta}{s\beta+1}}(\log D)^{\frac{s\beta-s}{1+s\beta}}$.
  • If $s<1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim 1$, $r_{\mathrm{opt}}\gtrsim D^{\frac{s\beta+1-\beta}{\beta-1}}\log D$ and $\mathcal{E}_{\mathrm{opt}}\eqsim D^{-s}$.
Compared with the exp‐decay LRS, both regimes enjoy an additional logarithmic improvement in excess‐risk decay. In particular, for the hard‐learning regime, the logarithmic factor is completely eliminated. To achieve such improvement, the decay-phase duration only needs to scale sublinearly with $D$, as indicated by $r_{\mathrm{opt}}\to 0$ as $D\to\infty$. This aligns with the WSD practice in LLM pre-training, where the decay phase typically occupies only 10%-20% of the total training duration. Moreover, our theory suggests that for harder tasks, the decay fraction can be reduced even further to enhance compute efficiency.
Compute-optimal scaling.
Analogous improvements hold in the compute-limited regime. Assuming $b=a/K$ and imposing the compute constraint $MD=C$, the compute-optimal satisfies:
  • If $s\geqslant 1-1/\beta$, then $\gamma_{\mathrm{opt}}\eqsim(C/\log C)^{-\frac{1+s\beta-\beta}{2+s\beta}},r_{\mathrm{opt}}\eqsim 1,M_{\mathrm{opt}}\eqsim(C/\log C)^{\frac{1}{2+s\beta}}$, $D_{\mathrm{opt}}\eqsim C^{\frac{1+s\beta}{2+s\beta}}(\log C)^{\frac{1}{2+s\beta}},$ and $\mathcal{E}_{\mathrm{opt}}\eqsim C^{-\frac{s\beta}{2+s\beta}}(\log C)^{\frac{s\beta-s}{2+s\beta}}.$
@@ 6 Experiments
6.1 Power-Law Kernel Regression
While the FSL is derived in the continuous-time limit, we now verify that it also accurately captures the loss dynamics and scaling behavior of the discrete-time SGD (4). Specifically, we consider the PLK regression with difficulty $s=0.5$ and capacity $\beta=4$, corresponding to a hard-learning regime and the results are shown in Figure 1.
FSL accurately captures the loss dynamics of SGD.
Figure 1(left) compares the loss dynamics of SGD with the predictions of the FSL under three representative LRSs: cosine, WSD, and an unconventional cyclical schedule [57]. Across all cases, the FSL provides a remarkably accurate description of the SGD’s loss evolution. Comparing the WSD and cosine schedules, we observe that the loss under WSD exhibits a slower decay during the stable phase but undergoes a much sharper drop once the decay phase begins, ultimately yielding a lower final loss. This seemingly counterintuitive two-phase dynamical behavior of WSD aligns well with empirical observations in practical LLM pre-training [22, 65, 59].
FSL predicts the scaling behavior of SGD.
Figure 1(right) further validates the scaling laws derived in Section 5 for the three canonical LRSs—constant, exponential, and WSD-like (25). The results show that the final-step loss of SGD closely follows the theoretical predictions of FSL. Among these schedules, WSD yields the best scaling performance, followed by exponential decay, while the constant schedule performs the worst.
More experiment details and additional results experiments with varying $(s,\beta)$ and other LRSs are provided in Appendix B.1, and exhibit consistent behaviors.
@@ 6.2 LLM Pre-training
We now evaluate the practical utility of FSL as a surrogate model for capturing the loss dynamics of LLM pre‑training. Specifically, three popular LRSs: cosine, WSD, and the 8‑1‑1 [5] are considered; see Figure 2(b)(left) for a visualization. In the 8‑1‑1 LRS, the learning rate is reduced by a factor $\sqrt{10}$ at 80% and 90% of the total token budget, yielding a final value that is $0.1$ times the peak learning rate. For more experiment details, we refer to Appendix B.2.
@@
FSL accurately fit and predict loss curves.
We first quantify the descriptive and predictive power of FSL. Following the protocol of [60] and [40], we restrict attention to the post‑warmup portion of the loss trajectory. Two Llama [61] models (400 M and 1 B) are trained on 20 B tokens under the three LRSs. For each model we (i) fit the FSL parameters on the loss curve obtained using the 8‑1‑1 LRS and (ii) deploy the fitted FSL to predict the loss curves of the cosine and WSD schedules. Figure 2(a) demonstrates that FSL not only fits the 8‑1‑1 trajectory accurately but also generalizes reliably to the unseen WSD and cosine schedules for both model sizes.
The FSL‑optimal LRS is WSD-like.
We next leverage the fitted FSL to design improved LRSs. Specifically, we numerically minimize the final-step loss over the space of LRSs using the fitted FSL. This experiment employs a 1B-parameter QwenMoE model [70], trained on 20B tokens using the same three LRSs. We fit the FSL using the trajectory from the 8-1-1 LRS and numerically solve for the FSL-optimal LRS. The model is then trained under this FSL-optimal LRS, using the same compute budget, and compared against the baseline LRSs.
Figure 2(b)(left) shows that surprisingly, the FSL-optimal LRS is WSD-like and the decay phase drives the learning rate far below the conventional $0.1\eta_{\max}$ threshold. This echos recent empirical recommendations by [4, 18, 40]. Furthermore, Figure 2(b)(right) demonstrates that the FSL‑optimal schedule yields a strictly lower final loss than all baselines, substantiating its practical relevance.
Taken together, these results suggest that FSL offers a faithful surrogate model for studying LLM training dynamics and holds promise as a principled tool for interpreting and designing LRSs in large-scale pre-training.